A canoeist rows downstream in 1.5 hours and back upstream in 3 hours. What is the rate of the canoe if the distance rowed is 12 miles in each direction?

Answer by mananth(16949)

speed of boat = xmph
speed of current = y mph
against current -------- 3.00 hours
with current ------ 1.50 hours

Distance = same= 12
Boat speed =x
current speed =y
t=d/r
12 / ( x - y )= 3.00
3 ( x - y ) = 12.00
3 x -3 y = 12 ....................1
12 / ( x + y )= 1.50
1.50 ( x + y ) = 12
1.50 x + 1.50 y = 12 ...............2
Multiply (1) by 1.00
Multiply (2) by 2.00
we get
3 x + -3 y = 12
3 x + 3 y = 24
6 x = 36
/ 6
x = 6 mph Boat speed